# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

### v3

@@ Appendix A Appendix
A.1 Robustness Tests
@@
$IV(\hat{Dis})$0.3336*** -0.0479*** -0.0408** -0.0424**
(-602.14) (-3.0748) (-2.3966) (-2.3533)
$TobinQ$0.0007 -0.0027 -0.0013 -0.0004
(-0.5971) (-0.4037) (-0.2634) (-0.0756)
$Top10$ 3.476e-05 -0.0010 -0.0009 -0.0005
(0.3084) (-0.9808) (-0.9708) (-0.5052)
$ListAge$0.0056 0.0910** 0.0799** 0.0061
(-0.9005) (2.4981) (2.3788) (0.1734)
$Lev$0.0029 -0.1554 -0.1407 -0.1333
(-0.3495) (-1.3058) (-1.2618) (-1.1844)
$Size$0.0045** 0.3760*** 0.2594*** 0.2260***
(-2.6017) (14.301) (13.453) (10.029)
$Avg$0.0017 0.2278*** 0.2680** 0.2660**
(-0.7314) (2.6949) (2.7600) (2.6811)
$Der$ 0.0037** -$0$0193 -0.0216
(2.8094) (-1.9412) (-1.7701) (-1.9119)
$N.TFP_{province}$ 0.0100 0.0999 0.0851 0.0408
(0.5583) (0.6350) (0.5902) (0.2565)
@@
$F$ 2.983e+04 227.22 208.52 170.30
$N$ 16907 12892 12892 12892
R2 0.9542 0.1794 0.1671 0.1408
  • Note: The second stage of 2SLS regression follows the baseline regression with a one-period lag.
@@
A.2 Optuna Parameter Range & Correlation Coefficient
Table 12: Correlation Matrix

### v4

@@ Appendix A Appendix
A.1 Descriptive Statistics
Table 9: Descriptive Statistics of Main Variables
Variables Count Mean SD Min 25% 50% 75% Max
$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
$Master$ 10461 4.31 1.55 1.10 3.18 4.20 5.31 8.42
$Undergrad$ 12595 6.31 1.35 3.56 5.35 6.17 7.15 10.05
A.2 Correlation Coefficient
Table 10: Correlation Matrix
A.3 Robustness Tests and Endogeneity Treatment
Second, since the TFP of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al. (2024) as a control variable (see table 12).
Third, to further mitigate everse causality endogeneity issues, following Yang et al. (2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable. A two-stage least squares (2SLS) estimation is conducted, and the Kleibergen-Paap rk LM statistics verify the verify the validity of instrumental variables(see table 13).
Fourth, to address sample selection bias, we denote disagreements under 20% as non-existent and the remainder as present. Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table 14).
@@ A.3 Robustness Tests and Endogeneity Treatment
$IV(\hat{Dis})$0.3302*** -0.0476*** -0.0409** -0.0425**
(-328.74) (-3.0292) (-2.3843) (-2.3456)
$TobinQ$ 0.0002 -0.0027 -0.0013 -0.0003
(0.1447) (-0.3986) (-0.2574) (-0.0685)
$Top10$ 0.0001 -0.0009 -0.0009 -0.0005
(1.1390) (-0.9769) (-0.9679) (-0.5016)
$ListAge$0.0118 0.0906** 0.0797** 0.0058
(-1.2204) (2.4886) (2.3696) (0.1655)
$Lev$0.0061 -0.1555 -0.1408 -0.1335
(-0.7749) (-1.3069) (-1.2627) (-1.1853)
$Size$0.0044** 0.3761*** 0.2594*** 0.2260***
(-2.0583) (14.306) (13.451) (10.030)
$Avg$0.0015 0.2277*** 0.2680** 0.2660**
(-0.5293) (2.6949) (2.7600) (2.6812)
$Der$ 0.0045** -$0$0193 -0.0216
(3.1346) (-1.9398) (-1.7686) (-1.9099)
$N.TFP_{province}$0.0030 0.0994 0.0847 0.0404
(-0.2042) (0.6320) (0.5877) (0.2540)
@@
$F$ 2.049e+04 227.20 208.52 170.30
$N$ 12892 12892 12892 12892
Kleibergen-Paap rk LM 622.33****
R2 0.9517 0.1794 0.1671 0.1408
  • Note: TFP is t+1 period, just like the baseline regression setup.
@@
A.4 Optuna Parameter Range